please help me !! I am trying to help my son with an algebra word problem and I do not remember how to do this type of algebraic equation. I do remember: Time = Distance divided by Rate . We have filled in the variables but that is as far as we got. Thanks, here is the question.....
Fastball thrown by.....travels at a speed of 145km/hr toward the batter who is 18.3 meters away. How long does it take the baseball to travel this distance.
Distance divided by Rate = Time
18.3m 145km/hr = t

Found 2 solutions by checkley77, stanbon:
Answer by checkley77(12569)

D=RT or T=D/R
T=18.3 /145,000 (THERE ARE 1,000 METERS IN A KM)
T=.0001262 HRS. or .007572 min. or .4543 SEC.

Answer by stanbon(57953)

Fastball thrown by.....travels at a speed of 145km/hr toward the batter who is 18.3 meters away. How long does it take the baseball to travel this distance.
----------------------------------------
Distance divided by Rate = Time
145 km/hr = 145000 meters per hour
----------------------------------------
Equation:
18.3m / 145000 m/hr = 0.0001262...hrs
----
Change to seconds:
0.0001262 hrs = 0.0001262*3600 seconds = 0.4543 seconds
============================================================
Cheers,
Stan H.